OpenStudy (anonymous):

10. For y = 4x + 2x2 – x3, show how to find the value for y, when x = 2. Show your work Can some one explain how to solve this question with out giving me the answer please.

OpenStudy (daniellelovee):

replace the value where the x is on the equation

OpenStudy (daniellelovee):

y = 4(2) + 2x2 – (2)3

OpenStudy (anonymous):

Remember pemdas though which means that you have to do multiplication before adding and subtracting

OpenStudy (anonymous):

y=4(2)+2x2 4x2=8 2x2=4 Is that right

OpenStudy (anonymous):

yes then you do the adding and subtracting

OpenStudy (anonymous):

So do i add the 8 & 4 together

OpenStudy (anonymous):

yup then subtract by what?

OpenStudy (anonymous):

x^3

OpenStudy (anonymous):

8+4=12

OpenStudy (anonymous):

yes and its not x to the power of three but rather x times three

OpenStudy (anonymous):

so i take away 12 from x

OpenStudy (anonymous):

no because if you remember pemdas that stands for parenthesis, exponents, multiplication,devision, add subtract so you must do thing in an equation in that order so instead of taking 12 from x you would times x by three and then subtract 12 by that number to give you your answer

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

so it would be 12-3=9

OpenStudy (anonymous):

no you have to multiply 3 time x or 2 in this equation so what is 3x2

OpenStudy (anonymous):

6

OpenStudy (anonymous):

yup so what is 12-6?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

so what is Y then?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

exactly that is right

OpenStudy (anonymous):

so thats all i do right

OpenStudy (anonymous):

yup just remember pemdas and substitute the numer they give you for the variable or letter and do it in order to find the answer

OpenStudy (anonymous):

oh ok ty